Solve the system by substitution. 3x – y = –7, x + y = –9
Answer by jgr45(31)

Solve the system by substitution. 3x – y = –7, x + y = –9
To solve by substitution you need to isolate one of the variables by itself. Here either one works so I'll pick y.
y = -9 - x
Now, rewrite the first equation and solve it, replacing y with -9-x.
3x - (-9-x) = -7
3x + 9 + x = -7
4x + 9 = -7
4x = -16
x = -4
Plug -4 in for x:
y = -9 - (-4)
y = -5
So the point of intersection is (-4,-5).